# Video Understanding with Large Language Models: A Survey

### 29 Dec 2023 v1

@@ Abstract
With the burgeoning growth of online video platforms and the escalating volume of video content, the demand for proficient video understanding tools has intensified markedly. With Large Language Models (LLMs) showcasing remarkable capabilities in key language tasks, this survey provides a detailed overview of the recent advancements in video understanding harnessing the power of LLMs (Vid-LLMs). The emergent capabilities of Vid-LLMs are surprisingly advanced, particularly their ability for open-ended spatial-temporal reasoning combined with commonsense knowledge, suggesting a promising path for future video understanding. We examine the unique characteristics and capabilities of Vid-LLMs, categorizing the approaches into four main types: LLM-based Video Agents, Vid-LLMs Pretraining, Vid-LLMs Instruction Tuning, and Hybrid Methods. Furthermore, this survey also presents a comprehensive study of the tasks and datasets for Vid-LLMs, along with the methodologies employed for evaluation. Additionally, the survey explores the expansive applications of Vid-LLMs across various domains, thereby showcasing their remarkable scalability and versatility in addressing challenges in real-world video understanding. Finally, the survey summarizes the limitations of existing Vid-LLMs and the directions for future research. For more information, we recommend readers visit the repository at https://github.com/yunlong10/Awesome-LLMs-for-Video-Understanding.
@@
11footnotetext: *Equal contribution.22footnotetext: †Corresponding author.
@@ 1 Introduction
We live in a multimodal world where video has become the most common form of media, thanks in part to the advancement of Internet technology, especially mobile Internet technology. With the rapid expansion of online video platforms and the increasing popularity of cameras in surveillance, entertainment, and automatic driving, video content has emerged as a crucial and engaging medium, surpassing conventional forms of text and combined image-text in terms of richness and appeal. This advancement has promoted an exponential surge in video production, with millions of videos being created daily. However, manually processing this sheer volume of video content is laborious and time-consuming. Consequently, there is a growing demand for tools that can effectively manage, analyze, and process this abundance of video content. To meet this demand, video understanding and analysis technology has emerged, leveraging intelligent analysis techniques. This technology aims to automatically recognize and interpret video content, thereby alleviating the burden on human operators.
Figure 1: The development of video understanding methods can be summarized into four stages: (1) Conventional Methods, (2) Neural Video Models, (3) Self-supervised Video Pretraining, and (4) Large Language Models for Video Understanding, i.e., Vid-LLMs.
@@
Conventional Methods. In the early stages of video understanding, handcrafted feature extraction techniques such as Scale-Invariant Feature Transform (SIFT) (lindeberg2012scale, ), Speeded-Up Robust Features (SURF) bay2008speeded , and Histogram of Oriented Gradients (HOG) dalal2005histograms were used to capture key information in videos. Background subtraction sobral2014comprehensive , optical flow methods tu2022optical , and Improved Dense Trajectories (IDT) wang2013action ; shu2015action were used to model the motion information for tracking. Since videos can be viewed as time-series data, time-series analysis techniques like Hidden Markov Models (HMM) liu2003video were also used to understand video content. Before the popularity of deep learning, basic machine learning algorithms such as Support Vector Machines (SVM) sidenbladh2004detecting , Decision Trees, and Random Forests were also used in video classification and recognition tasks. Cluster analysis chan2008modeling for classifying video segments, or Principal Component Analysis (PCA) bouwmans2014robust ; hazelhoff2008video for data dimensionality reduction were also commonly used methods for video analysis.
Neural Video Models. Compared with classical methods, deep learning methods for video understanding possess superior task-solving capabilities.
@@
DeepVideo KarpathyCVPR14 was the earliest method introducing a deep neural network, specifically a Convolutional Neural Network (CNN), for video understanding. However, the performance was not superior to the best handcrafted-feature method due to the inadequate use of motion information. Two-stream networks feichtenhofer2016convolutional combined both CNN and IDT to capture the motion information to improve the performance, which verified the capability of deep neural networks for video understanding. To handle long-form video understanding, Long Short-Term Memory (LSTM) was adopted yue2015beyond . Temporal Segment Network (TSN) wang2016temporal was also designed for long-form video understanding by analyzing video segments individually and then aggregating them. Based on TSN, Fisher Vectors (FV) encoding sekma2015human , Bi-linear encoding diba2017deep , and Vector of Locally Aggregated Descriptors (VLAD) mironicua2016modified encoding were introduced li2022transvlad . These methods improved performance on the UCF-101 soomro2012ucf101 and HMDB51 kuehne2011hmdb datasets. Different from Two-stream networks, 3D networks started another branch by introducing 3D CNN to video understanding (C3D) tran2015learning . Inflated 3D ConvNets (I3D) carreira2017quo utilizes the initialization and the architecture of 2D CNN, Inception szegedy2015going , to gain a huge improvement on the UCF-101 and HMDB51 datasets. Subsequently, people began employing the Kinetics-400 (K-400) kay2017kinetics and Something-Something goyal2017something datasets to evaluate the model’s performance in more challenging scenarios. ResNet he2016deep , ResNeXt Xie_2017_CVPR , and SENet Hu_2018_CVPR were also adapted from 2D to 3D, resulting in the emergence of R3D Hara_2017_ICCV , MFNet Chen_2018_ECCV , and STC Diba_2018_ECCV . To improve the efficiency, the 3D network has been decomposed into 2D and 1D networks in various studies (e.g., S3D S3D , ECO zolfaghari2018eco , P3D qiu2017learning ). LTC LTC , T3D T3D , Non-local non-local , and V4D zhang2019v4d focus on long-form temporal modeling, while CSN tran2019video , SlowFast feichtenhofer2019slowfast , and X3D feichtenhofer2020x3d tend to attain high efficiency. The introduction of Vision Transformers (ViT) dosovitskiy2021an promotes a series of prominent models (e.g., TimeSformer bertasius2021space , VidTr li2021vidtr , ViViT arnab2021vivit , MViT fan2021multiscale ).
@@
Self-supervised Video Pretraining. Transferability li2023cross ; wang2023feature in self-supervised pretraining modelszhang2023rethinking for video understanding allows them to generalize across diverse tasks with minimal additional labeling, overcoming the early deep learning models’ need for extensive task-specific data. VideoBERT sun2019videobert is an early attempt to perform video pre-training. Based on the bidirectional language model BERT kenton2019bert , pertaining tasks are designed for self-supervised learning from video-text data. It tokenizes video features with hierarchical k-means. The pre-trained model can be fine-tuned to handle multiple downstream tasks, including action classification and video captioning. Following the "pre-training" and "fine-tuning" paradigm, a large number of studies on pre-trained models for video understanding, especially video-language models, have emerged. They either use different architectures (ActBERT zhu2020actbert , Masked Autoencoders as Spatiotemporal Learners feichtenhofer2022masked , OmniMAE girdhar2023omnimae , VideoMAE tong2022videomae , MotionMAE yang2022self ) or pre-training and fine-tuning strategies (MaskFeat wei2022masked , VLM:task-agnostic xu2021vlm , ALPRO li2022align , all-in-one transformer moritz2020all , maskViT gupta2022maskvit , CLIP-ViP xue2022clip , Revealing Single Frame Bias for Video-and-Language Learning lei2022revealing , LF-VILA sun2022long , EMCL jin2022expectation , HiTeA ye2023hitea , CHAMPAGNE han2023champagne ).
@@
This survey is structured as follows: Section 2 offers a comprehensive overview, emphasizing methods that harness the capabilities of LLMs and detailing the specific tasks and datasets these methods address. In Section 3, we delve into details of recent researches leveraging LLMs for video understanding, presenting their unique approaches and impact in the field. Section 4 offers a detailed summary and analysis of various tasks, their associated datasets, and evaluation metrics. Section 5 explores the application of Video-LLMs across multiple significant fields. The survey concludes in Section 6 by summarizing key findings and identifying unresolved challenges and potential areas for future research.
@@ 2 Foundations
Video understanding is a challenging task that has inspired the creation of numerous innovative tasks aimed at augmenting a model’s capability to interpret video content. From the foundational task of video classification and action recognition, which categorizes videos into predefined classes, the field has evolved to include more intricate tasks. These range from captioning videos with detailed descriptions, to video question answering. The latter not only demands an understanding of the video content but also requires reasoning about logical and commonsense knowledge to formulate responses. As we advance in this field, tasks are becoming increasingly complex and challenging, necessitating models that can interpret videos as intuitively as humans do. We summarize the main tasks in video understanding as follows:
Recognition and Anticipation. These tasks form a cohesive duo in video understanding, emphasizing the temporal continuity and progression in videos.
Captioning and Summarization. Focusing on more fine-grained details, these tasks involve providing accurate and specific textual descriptions of each moment and distilling the essence of the video, capturing the overarching themes and key narratives. These tasks offer both a micro and macro understanding of video content, combining detail-oriented insights with a broader perspective.
@@
2.1 Vision Integration with LLM
@@
2.2 Role of Language in Video Understanding
@@ 3.3 Vid-LLM Instruction Tuning
In general, fine-tuning pre-trained large models is compute-intensive. Due to limited computational resources, not all parameters of the model are updated during fine-tuning of large models; instead, some parameters of adapters are updated. In Vid-LLMs instruction tuning, two types of adapters, connective and insertive, are commonly used. Some methods also mix these two types of adapters. The connective adapter is usually placed between the visual backbone and the LLM, mainly used for aligning visual and textual semantics, while the insertive adapter is typically inserted into the LLM.
@@ 4.5.1 Pretraining Datasets
4.5.2 Fine-tuning Datasets

### 04 Jan 2024 v2

@@ Abstract
With the burgeoning growth of online video platforms and the escalating volume of video content, the demand for proficient video understanding tools has intensified markedly. Given the remarkable capabilities of Large Language Models (LLMs) in language and multimodal tasks, this survey provides a detailed overview of the recent advancements in video understanding harnessing the power of LLMs (Vid-LLMs). The emergent capabilities of Vid-LLMs are surprisingly advanced, particularly their ability for open-ended spatial-temporal reasoning combined with commonsense knowledge, suggesting a promising path for future video understanding. We examine the unique characteristics and capabilities of Vid-LLMs, categorizing the approaches into four main types: LLM-based Video Agents, Vid-LLMs Pretraining, Vid-LLMs Instruction Tuning, and Hybrid Methods. Furthermore, this survey presents a comprehensive study of the tasks, datasets, and evaluation methodologies for Vid-LLMs. Additionally, it explores the expansive applications of Vid-LLMs across various domains, highlighting their remarkable scalability and versatility in real-world video understanding challenges. Finally, it summarizes the limitations of existing Vid-LLMs and outlines directions for future research. For more information, readers are recommended to visit the repository at https://github.com/yunlong10/Awesome-LLMs-for-Video-Understanding.
@@
11footnotetext: Equal contribution.22footnotetext: Corresponding author.
@@ 1 Introduction
We live in a multimodal world where video has become the predominant form of media. With the rapid expansion of online video platforms and the growing prevalence of cameras in surveillance, entertainment, and autonomous driving, video content has risen to prominence as a highly engaging and rich medium, outshining traditional text and image-text combinations in both depth and appeal. This advancement has promoted an exponential surge in video production, with millions of videos being created daily. However, manually processing such a sheer volume of video content proves to be both labor-intensive and time-consuming. Consequently, there is a growing demand for tools that can effectively manage, analyze, and process this abundance of video content. To meet this demand, video understanding methods and analysis technologies have emerged, leveraging intelligent analysis techniques to automatically recognize and interpret video content, thus significantly reducing the workload on human operators. Moreover, the continuous advancements in these methods are enhancing their task-solving capabilities, enabling them to handle a diverse range of video understanding tasks with increasing proficiency.
Figure 1: The development of video understanding methods can be summarized into four stages: (1) Conventional Methods, (2) Neural Video Models, (3) Self-supervised Video Pretraining, and (4) Large Language Models for Video Understanding, i.e., Vid-LLMs. Their task-solving capability is continuously improving, and they possess the potential for further enhancement.
@@
Conventional Methods. In the early stages of video understanding, handcrafted feature extraction techniques such as Scale-Invariant Feature Transform (SIFT) (lindeberg2012scale, ), Speeded-Up Robust Features (SURF) bay2008speeded , and Histogram of Oriented Gradients (HOG) dalal2005histograms were used to capture key information in videos. Background subtraction sobral2014comprehensive , optical flow methods tu2022optical , and Improved Dense Trajectories (IDT) wang2013action ; shu2015action were used to model the motion information for tracking. Since videos can be viewed as time-series data, time-series analysis techniques like Hidden Markov Models (HMM) liu2003video were also used to understand video content. Before the popularity of deep learning, basic machine learning algorithms such as Support Vector Machines (SVM) sidenbladh2004detecting , Decision Trees yuan2002automatic , and Random Forests were also used in video classification and recognition tasks. Cluster analysis chan2008modeling for classifying video segments, or Principal Component Analysis (PCA) bouwmans2014robust ; hazelhoff2008video for data dimensionality reduction were also commonly used methods for video analysis.
@@
Neural Video Models. Compared with classical methods, deep learning methods for video understanding possess superior task-solving capabilities. DeepVideo KarpathyCVPR14 was the earliest method introducing a deep neural network, specifically a Convolutional Neural Network (CNN), for video understanding. However, the performance was not superior to the best handcrafted-feature method due to the inadequate use of motion information. Two-stream networks feichtenhofer2016convolutional combined both CNN and IDT to capture the motion information to improve the performance, which verified the capability of deep neural networks for video understanding. To handle long-form video understanding, Long Short-Term Memory (LSTM) was adopted yue2015beyond . Temporal Segment Network (TSN) wang2016temporal was also designed for long-form video understanding by analyzing video segments individually and then aggregating them. Based on TSN, Fisher Vectors (FV) encoding sekma2015human , Bi-linear encoding diba2017deep , and Vector of Locally Aggregated Descriptors (VLAD) mironicua2016modified encoding were introduced li2022transvlad . These methods improved performance on the UCF-101 soomro2012ucf101 and HMDB51 kuehne2011hmdb datasets. Different from Two-stream networks, 3D networks started another branch by introducing 3D CNN to video understanding (C3D) tran2015learning . Inflated 3D ConvNets (I3D) carreira2017quo utilizes the initialization and the architecture of 2D CNN, Inception szegedy2015going , to gain a huge improvement on the UCF-101 and HMDB51 datasets. Subsequently, people began employing the Kinetics-400 (K-400) kay2017kinetics and Something-Something goyal2017something datasets to evaluate the model’s performance in more challenging scenarios. ResNet he2016deep , ResNeXt Xie_2017_CVPR , and SENet Hu_2018_CVPR were also adapted from 2D to 3D, resulting in the emergence of R3D Hara_2017_ICCV , MFNet Chen_2018_ECCV , and STC Diba_2018_ECCV . To improve the efficiency, the 3D network has been decomposed into 2D and 1D networks in various studies (e.g., S3D S3D , ECO zolfaghari2018eco , P3D qiu2017learning ). LTC LTC , T3D T3D , Non-local non-local , and V4D zhang2019v4d focus on long-form temporal modeling, while CSN tran2019video , SlowFast feichtenhofer2019slowfast , and X3D feichtenhofer2020x3d tend to attain high efficiency. The introduction of Vision Transformers (ViT) dosovitskiy2021an promotes a series of prominent models (e.g., TimeSformer bertasius2021space , VidTr li2021vidtr , ViViT arnab2021vivit , MViT fan2021multiscale ).
@@
Self-supervised Video Pretraining. Transferability li2023cross ; wang2023feature in self-supervised pretraining modelszhang2023rethinking for video understanding allows them to generalize across diverse tasks with minimal additional labeling, overcoming the early deep learning models’ need for extensive task-specific data. VideoBERT sun2019videobert is an early attempt to perform video pre-training. Based on the bidirectional language model BERT kenton2019bert , pertaining tasks are designed for self-supervised learning from video-text data. It tokenizes video features with hierarchical k-means. The pre-trained model can be fine-tuned to handle multiple downstream tasks, including action classification and video captioning. Following the "pre-training" and "fine-tuning" paradigm, a large number of studies on pre-trained models for video understanding, especially video-language models, have emerged. They either use different architectures (ActBERT zhu2020actbert , SpatiotemporalMAE feichtenhofer2022masked , OmniMAE girdhar2023omnimae , VideoMAE tong2022videomae , MotionMAE yang2022self ) or pre-training and fine-tuning strategies (MaskFeat wei2022masked , VLM xu2021vlm , ALPRO li2022align , All-in-One transformer moritz2020all , maskViT gupta2022maskvit , CLIP-ViP xue2022clip , Singularity lei2022revealing , LF-VILA sun2022long , EMCL jin2022expectation , HiTeA ye2023hitea , CHAMPAGNE han2023champagne ).
@@
This survey is structured as follows: Section 2 offers a comprehensive overview, emphasizing methods that harness the capabilities of LLMs and detailing the specific tasks and datasets these methods address. In Section 3, we delve into details of recent research leveraging LLMs for video understanding, presenting their unique approaches and impact in the field. Section 4 offers a detailed summary and analysis of various tasks, their associated datasets, and evaluation metrics. Section 5 explores the application of Video-LLMs across multiple significant fields. The survey concludes in Section 6 by summarizing key findings and identifying unresolved challenges and potential areas for future research.
@@ 2 Foundations
Video understanding is a challenge that has inspired the creation of numerous innovative tasks to enhance models’ video interpretation capabilities. The field has evolved from foundational tasks like video classification and action recognition, which categorize videos into predefined classes, to encompass more intricate tasks. These include captioning videos with detailed descriptions and answering questions about video content. The latter not only requires an understanding of the video but also involves reasoning with logical and commonsense knowledge. As the field advances, tasks are becoming increasingly complex and challenging, calling for models with near-human levels of video interpretation. We summarize the main tasks in video understanding as follows:
Recognition and Anticipation. These tasks are closely linked in video understanding, focusing on the temporal continuity and progression within videos.
Captioning and Summarization. Focusing on more fine-grained details, these tasks involve creating accurate and specific textual descriptions for each moment and distilling the video’s essence, thereby encapsulating its overarching themes and key narratives. They provide both detailed and broad perspectives of the video content.
@@
2.1 Vision Integration with LLMs
@@
2.2 Language’s Roles in Video Understanding
@@ 3.3 Vid-LLM Instruction Tuning
In general, fine-tuning pre-trained large models is compute-intensive. Due to limited computational resources, not all parameters of the model are updated during fine-tuning of large models hu2021lora ; hua2021noise ; hua2022fine ; instead, some parameters of adapters are updated. In Vid-LLMs instruction tuning, two types of adapters, connective and insertive, are commonly used. Some methods also mix these two types of adapters. The connective adapter is usually placed between the visual backbone and the LLM, mainly used for aligning visual and textual semantics, while the insertive adapter is typically inserted into the LLM.
@@ 4.5.1 Pretraining Dataset
4.5.2 Fine-tuning Dataset